OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = x2 + 3x – 10? x–intercepts: (–2, 0) and (5, 0) y–intercept: (0, 10) x–intercepts: (–2, 0) and (–5, 0) y–intercept: (0, 10) x–intercepts: (–2, 0) and (5, 0) y–intercept: (0, –10) x–intercepts: (2, 0) and (–5, 0) y–intercept: (0, –10)

OpenStudy (anonymous):

y = x2 + 3x – 10? for y intercept, put x=0 y = 0 + 0 - 10 y = -10 <------- y intercept for x intercept, put y=0 0 = x^2 + 3x - 10 0 = x^2 + 5x - 2x - 10 0 = x(x + 5) -2(x + 5) 0 = (x - 2)(x + 5) so, x = 2 or x = -5 x–intercepts: (2, 0) and (–5, 0) y–intercept: (0, –10)